OpenStudy (anonymous):

how does integral of 4dx/(sqrt(4-x^2) turn into 4arcsin(x/2)?

zepdrix (zepdrix):

Are you familiar with Trigonometric Substitutions?

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{4dx}{\sqrt{4-x^2}}\]

OpenStudy (anonymous):

no i'm not

zepdrix (zepdrix):

Hmm

OpenStudy (anonymous):

i understand that 1/sqrt(1-x^2 turns into arcsin

zepdrix (zepdrix):

I guess another option would be to do a little algebra, make a u-substitution, and then relate it to the integral which produces arcsine from a table or something.

zepdrix (zepdrix):

You do? Ok great. Let's turn it into that then :)

OpenStudy (anonymous):

ok!

zepdrix (zepdrix):

Inside the root, let's pull a 4 out of each term,\[\Large\rm \int\limits\limits \frac{4dx}{\sqrt{4\left(1-\frac{x^2}{4}\right)}}\]This is going to get a little messy, so try to follow along. We want to bring the 4 into the square with the x. 4 = 2^2, yes? So it comes in as a 2.\[\Large\rm \int\limits\limits\limits \frac{4dx}{\sqrt{4\left(1-\left[\frac{x}{2}\right]^2\right)}}\]

OpenStudy (anonymous):

hmm ok

zepdrix (zepdrix):

Let's pull the 4 out of the root as a 2, and simplify the 2 on bottom and 4 on top.\[\Large\rm 2\int\limits \frac{dx}{\sqrt{1-\left[\frac{x}{2}\right]^2}}\]

OpenStudy (anonymous):

ok got that

zepdrix (zepdrix):

We are trying to get the form that you mentioned. So we'll make the substitution \(\Large\rm u=\dfrac{x}{2}\)

zepdrix (zepdrix):

What do you get for the differential? \(\Large\rm du=?\)

OpenStudy (anonymous):

ahhh interesting!

OpenStudy (anonymous):

you get 1/2

zepdrix (zepdrix):

\[\Large\rm du=\frac{1}{2}dx\]Ok great!

zepdrix (zepdrix):

Oh let's multiply the 1/2 to the other side,\[\Large\rm 2du=dx\]And then just plug everything in. Should work out really nicely from here.

OpenStudy (anonymous):

so you just multiply another 2 outside

zepdrix (zepdrix):

\[\Large\rm 2\int\limits\limits \frac{2du}{\sqrt{1-\left[u\right]^2}}\]Yes, pull the new 2 outside with the other one, giving us a 4.

OpenStudy (anonymous):

Thank you so much!

zepdrix (zepdrix):

As a side note, we could have initially made the substitution \(\Large\rm x=2u\). That would have avoided some of the algebra sets. I wasn't quite sure which method would make more sense to you though. No problem! :)

zepdrix (zepdrix):

algebra steps*

OpenStudy (anonymous):

how do u make the substitution of x=2u?

zepdrix (zepdrix):

The reason to do it that way is to try and match the 4's up. \(\Large\rm x=2u\) Plugging it in gives us:\[\Large\rm \int\limits\limits \frac{4dx}{\sqrt{4-(2u)^2}}=\int\limits\frac{4dx}{\sqrt{4-4u^2}}\]And from there you can factor out the 4 and get that nice form that we got using the other method. You still have to replace the dx, but maybe this way is a little simpler not pulling the 4 into the square and all that. I dunno, you can decide for yourself which way makes more sense.

OpenStudy (anonymous):

that is interesting as well, thank you for your tips!